The width of a rectangular rug is one foot smaller than its length. The area of the rug is 30 square feet. Find the length and the width. Whats this in feet???
Answer by mathworm7271(5)

Let's say x is the length
then x-1 is the width
We know that Area=Length(width)
So, we substitute...
30=x(x-1)
Use the Distributive property
30=x^2-x
Subtract 30 from each side
0=x^2-x-30
Now we have to factor the Trinomial
0=(x-6)(x+5)
So that means either x-6=0 or x+5=0
Distribute and x=6 or x=-5
Since the length of a rug cannot be negative, x=6
So, the length is 6 and the width is 6-1 which is 5.
Make sure you have the proper units (feet)